OpenStudy (anonymous):

Subtract \[\frac{ 5 }{ x-2 } - \frac{ 3 }{ x-2 }\]

OpenStudy (anonymous):

I think it comes out as \[\frac{ 2 }{ (x+2)(x-2) }\]

OpenStudy (anonymous):

since the denominator is the same all you have to do is subtract the top numbers like you did and just leave the denominator the same

OpenStudy (anonymous):

but the answer choices look like this though http://awesomescreenshot.com/0ed319o185

OpenStudy (anonymous):

\(\Large \frac{5}{x-2}-\frac{3}{x-2}=\frac{2}{x-2}\) multiply \(\Large x+2\) with numerator and denominator of \(\Large \frac{2}{x-2}\) \(\Large \frac{2(x+2)}{(x-2)(x+2)}=\frac {2x+4}{(x-2)(x+2)} \)

OpenStudy (anonymous):

I see, thank you very much